# On TinyML and Cybersecurity: Electric Vehicle Charging Infrastructure Use Case

### v2

@@ VI-C Attack Scenario
In this work, an attack scenario proposed by ElHussini et al. is specifically focused on [101]. In this attack, an adversary could manipulate an EVCS to cause a traffic bottleneck by altering the charging schedules of EVs. The attack unfolds in several steps: First, the attacker locates the EVCS using tools such as Shodan and Censys, gaining access to IP addresses and ports. Then, by exploiting security weaknesses such as default usernames and passwords found in manufacturers’ manuals, the attacker takes control of the EVCS. This control enables the manipulation of EV charging schedules, redirecting charging traffic to specific locations and times, ultimately creating a bottleneck in the power grid. Finally, the attacker triggers disruptions at the grid level by coordinating sudden surges in power demand or reverse power flows from EVs back to the grid. These orchestrated attacks can lead to frequency instability, progressive failures, and imbalances between power demand and supply.
@@
There are several methods to counter the aforementioned attack. The first line of defense involves securing networks with firewalls and filtering connections from trusted sources, alongside implementing strong authentication protocols. These measures are designed to prevent common exploits and protect against basic attempts at unauthorized access. Following these initial steps, an additional layer of security can be introduced to monitor for unusual activity in how the EVs are scheduled to charge. Anomaly detection systems play a crucial role in this aspect, as they continuously monitor charging schedules, quickly identifying and responding to any deviations from normal patterns [101].
@@
TinyML emerges as an effective tool for implementing IDS within EVs. Considering that energy consumption and battery life are significant concerns in EVs, TinyML offers a viable solution. A key advantage of TinyML is its ability to conduct real-time anomaly detection directly on the microcontrollers embedded within EVs. Processing data locally on these devices significantly reduces the risk of data interception and unauthorized access. This aspect is crucial because it reduces the amount of sensitive data transmitted over the network, thereby enhancing the security of the electric vehicle charging infrastructure.
VI-D TinyML Deployment
@@
An ML model is aimed to be trained in a computationally robust environment, specifically using CMS, and then optimized and deployed for inference on an edge device, which refers to a simple microcontroller in EV. This approach draws from multiple methodologies, including but not limited to the Edge Learning Machine (ELM) framework, as one example, where models are trained on a desktop environment and subsequently utilized for predictions on STM microcontrollers [24].
@@
Integrating TinyML into the vehicle’s internal systems significantly enhances security. A microcontroller, designed for IDS purposes, is installed within the EV to monitor the CAN bus. This monitoring is crucial as the CAN bus is the pathway for all operational data, managing various functions [110], including interactions with an EVCS. By examining traffic and communications for anomalies, the system helps identify attacks that could bypass traditional security measures such as firewalls. It detects unusual or malicious activities, such as those from compromised charging stations or incorrect charging data, which could affect the vehicle’s performance. Deploying this capability directly within the vehicle ensures a comprehensive defense mechanism, creating a protective shield that complements external security measures provided by the EVCS.
@@
VII Experimental Setup and Results
@@
VII-A Dataset Description
The CICIDS2017 dataset, provided by the Canadian Institute for Cybersecurity, serves as a foundational resource for our study. This comprehensive dataset is particularly useful for cybersecurity research and has been crafted to include a diverse range of common cyber-attacks [111, 112]. Using this dataset, our IDS is designed to identify various types of attacks, specifically: Dos (encompassing DoS Hulk, DoS GoldenEye, DoS slowloris, and DoS Slowhttptest), PortScan, DDoS, Brute Force (including FTP-Patator and SSH-Patator), Web Attack (comprising Web Attack-Brute Force, Web Attack-XSS, and Web Attack-SQL Injection), and a combined category of Bot/Infiltration/Heartbleed [113].
VII-B EVCI Dataset Creation
@@ VII-G TinyML Robustness
VIII Practical Implementation and Results
@@
IX Future Research Directions
IX-A Decentralized Federated Learning Approaches
@@
IX-B Adapting to New Threats using Online Learning
@@
IX-C Integrating Explainable AI (XAI) with TinyML

### v3

@@ V-C Cyber Attack Scenarios in EVCI
In an attack scenario proposed by ElHussini et al. [103], an adversary could manipulate an EVCS to cause a traffic bottleneck by altering the charging schedules of EVs. The attack unfolds in several steps: First, the attacker locates the EVCS using tools such as Shodan and Censys, gaining access to IP addresses and ports. Then, by exploiting security weaknesses such as default usernames and passwords found in manufacturers’ manuals, the attacker takes control of the EVCS. This control enables the manipulation of EV charging schedules, redirecting charging traffic to specific locations and times, ultimately creating a bottleneck in the power grid. Finally, the attacker triggers disruptions at the grid level by coordinating sudden surges in power demand or reversing power flows from EVs back to the grid. These orchestrated attacks can lead to frequency instability, progressive failures, and imbalances between power demand and supply.
There are several other attacks, such as FDI, where attackers manipulate data transmission between EVSEs and management systems, leading to overcharging batteries or disrupting grid stability [110]. Malware injections target publicly accessible EVSE units, introducing malicious software that can compromise the entire EVCS ecosystem, including the CMS and the power grid [110].
Additionally, poor database implementation can be exploited to insert, update, or delete data in the database through SQL injection attacks. This vulnerability enables attackers to inject malicious SQL commands to manipulate the database. They can execute commands affecting users’ ability to charge, modify charging station location data, or alter station availability status, potentially causing public safety issues [112].
To mitigate these risks, it is crucial to implement comprehensive cybersecurity measures in EVCI. The STRIDE framework helps identify potential security risks and vulnerabilities within EVCSs early in development. The six threats outlined in STRIDE include Spoofing, where attackers impersonate authorized sources to install malware on chargers, potentially causing station shutdowns; Tampering, involving the injection of false charging data that can mislead grid operations or damage EVs; Repudiation, where compromised chargers deny processing a charging fee, even if the EV owner has already paid for it, causing a financial loss to the EV owner; Information Disclosure, exposing sensitive communication between chargers and vehicles to unauthorized interception; DoS attacks that can flood the network and disable charging services; and Elevation of Privilege, where attackers gain unauthorized access to critical data and modify system files to disrupt the entire charging network. By using STRIDE, developers can proactively identify and address these threats, improving the safety and reliability of EVCI [117].
V-D TinyML: A Key Countermeasure in Cybersecurity for EVCI
@@
There are several methods to counter the attacks in the EVCI. The first line of defense involves securing networks with firewalls and filtering connections from trusted sources, alongside implementing strong authentication protocols. These measures are designed to prevent common exploits and protect against basic attempts at unauthorized access. Following these initial steps, an additional layer of security can be introduced to monitor for unusual activity in how the EVs are scheduled to charge. IDSs play a crucial role in this aspect, as they continuously monitor data traffic and user interactions, such as charging schedules, quickly identifying and responding to any deviations from normal patterns [103].
IDSs utilize various approaches based on the source of information and analysis strategies to safeguard against cyber threats. Host-based IDSs are deployed on individual hosts, using system, application logs, and audit trails to monitor and respond to malicious activities directly affecting the system. These IDSs focus on detecting unauthorized access, file tampering, and resource misuse by analyzing data from firewalls, routers, and operating systems. Conversely, network-based IDSs monitor the entire network’s traffic to identify potential threats affecting multiple systems, relying on network sensors to manage the vast amounts of data for efficient threat detection [118].
Although network-based IDSs implemented at charging stations play a crucial role, this paper primarily focuses on enhancing security from within EVs themselves. Integrating host-based IDSs directly within EVs provides critical defense against specific threats at the point of interaction between the EV and the charging station, such as FDI and spoofing. By equipping EVs with robust host-based IDS, a comprehensive defense mechanism is established that not only protects the vehicle’s systems but also secures data interactions during the charging process. This approach is pivotal in maintaining the integrity and security of the EV within the broader ecosystem of EVCI.
@@
Considering that energy consumption and battery life are significant concerns in EVs, TinyML emerges as an effective tool for implementing IDS within these vehicles. A key advantage of TinyML is its ability to conduct real-time anomaly detection directly on the microcontrollers embedded within EVs. By processing data locally on these devices, the risk of data interception and unauthorized access is significantly reduced. This localized approach not only enhances security but also aligns with the sustainability goals of reducing overall energy usage in electric vehicles.
@@
For this purpose, an ML model is aimed to be trained in a computationally robust environment, specifically a CMS, and then optimized and deployed for inference on an edge device, which in this case refers to a simple microcontroller in an EV. This approach draws from multiple methodologies, including, but not limited to, the Edge Learning Machine (ELM) framework. In this framework, models are trained in a desktop environment and subsequently utilized for predictions on STM microcontrollers [24].
@@
The microcontroller, designed for IDS purposes, is installed within the EV to monitor the CAN bus. This monitoring is crucial as the CAN bus serves as the pathway for all operational data, managing various functions, including interactions with an EVCS [119]. By examining traffic and communications for anomalies, the system identifies potential attacks that could bypass traditional security measures such as firewalls. It detects unusual or malicious activities, including those from compromised charging stations or incorrect charging data, which could compromise the vehicle’s performance.
@@
V-E TinyML limitations in EVCI and potential mitigation strategies
Incorporating TinyML into EVCI presents a promising solution for enhancing security. However, the deployment of TinyML in this context is not without its challenges. This section explores the inherent limitations of integrating TinyML within EVCI, ranging from computational constraints to issues of scalability and interoperability. It also discusses potential mitigation strategies to overcome these barriers.
V-E1 Computational Power
Computational power in EVs is constrained due to the need for energy conservation [120]. Although TinyML models are optimized, implementing them in EVs remains a significant challenge. To mitigate this challenge, an effective strategy is to use model optimization. Different pruning and quantization techniques [121] can reduce model size and computational requirements, improving the efficiency of ML models on resource-constrained devices.
V-E2 Real-Time Processing
The real-time nature of IoT-based EVCSs demands instantaneous data flow for efficient and secure charging operations [122]. To make TinyML real-time, techniques such as quantization-aware training (QAT), pruning, and post-training quantization (PTQ) can significantly reduce model size and computational requirements while maintaining high accuracy, thereby reducing delay.
V-E3 Memory Capacity
EVs have limited storage space, which restricts the deployment of larger models. One strategy to mitigate this issue is to use memory-efficient model architectures and techniques such as feature selection to minimize data input size.
V-E4 Implementation Limitations
Since microcontrollers equipped with TinyML models are integrated into the EV’s architecture, they can only be implemented in newly manufactured vehicles and not in older models. As a result, older vehicles lack the necessary hardware infrastructure to
V-E5 Threat Adaptability
Scalability and adaptability to evolving threats remain challenging due to the dynamic and multifaceted nature of the IoT-based environment of EVCI [122]. Continuous learning [123] is an effective technique to update TinyML models based on new attack data so that anomaly detection models can identify previously unseen threats.
V-E6 Standardization and Interoperability
The diversity in communication protocols used by charging stations (e.g., SAE, IEC 61851, ISO 15118, CHAdeMO) poses a challenge for developing a universal TinyML model. One strategy to address this issue is to design modular models that can adapt to different protocols through customizable preprocessing and feature extraction layers.
VI Experimental Setup and Results
@@
VI-A Dataset Description
The EVCI is vulnerable to a wide range of cyberattacks, making robust anomaly detection crucial. For this purpose, the CICIDS2017 dataset [124, 125] is selected, which, despite containing general attacks, aligns well with the types of threats found in EVCI. Attacks such as DoS or web attacks, which are included in the CICIDS2017 dataset, are also prevalent in electric vehicle charging infrastructures, as discussed in previous sections.
The CICIDS2017 dataset is provided by the Canadian Institute for Cybersecurity and serves as a foundational resource for cybersecurity research. In more detail, this dataset includes a diverse range of common cyber-attacks, including DoS (encompassing DoS Hulk, DoS GoldenEye, DoS slowloris, and DoS Slowhttptest), PortScan, DDoS, Brute Force (including FTP-Patator and SSH-Patator), Web Attack (comprising Web Attack-Brute Force, Web Attack-XSS, and Web Attack-SQL Injection), and a combined category of Bot/Infiltration/Heartbleed [126].
Using IoT-based datasets for anomaly detection in EVCI has been demonstrated in other studies as well. For instance, in a paper by Kilichev et al.[122], EVCSs are safeguarded using robust IDS. The authors employ a deep learning-based model evaluated on the Edge-IIoTset dataset[127], a comprehensive real-world IoT security dataset, to enhance detection accuracy and optimize resource efficiency.
Overall, the CICIDS2017 dataset provides a set of realistic traffic scenarios, enabling comprehensive detection capabilities that improve the reliability, safety, and efficiency of EVCI, therefore selected for this research.
VI-B EVCI Dataset Creation
@@ VI-G TinyML Robustness
VII Practical Implementation and Results
@@
VIII Future Research Directions
VIII-A Decentralized Federated Learning Approaches
@@
VIII-B Adapting to New Threats using Online Learning
@@
VIII-C Integrating Explainable AI (XAI) with TinyML